OpenStudy (anonymous):

Find the slope of the straight line that passes through the given points or state that the slope is undefined. Then indicate if the line through the points rises (left to right), falls (left to right), is horizontal, or is vertical 1b. (7, 1) and (7, 2)

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Plug them into the slope formula. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{2-1}{7-7}\)

OpenStudy (igreen):

Simplify 2 - 1 = ? 7 - 7 = ?

OpenStudy (anonymous):

2-1 = 1 7-7=0

OpenStudy (anonymous):

A vertical line passing through both points.

OpenStudy (anonymous):

Thank you guys! Can I ask another one with different points?

OpenStudy (igreen):

Yep, and anything divided by 0 is undefined..so the slope is undefined.

OpenStudy (igreen):

Sure, just close this and open a new one.

OpenStudy (anonymous):

Ohh okay cool, thank you. An I will